# SciHorizon: Benchmarking AI-for-Science Readiness from Scientific Data to Large Language Models

### v1

@@ 4.2 Results of LLM Assessment
Table 6: Performance of LLMs on Reasoning assessment across multiple scientific disciplines.
@@
Table 7: Performance of LLMs on Multimodality assessment across multiple scientific disciplines.
@@
Claude-3.5-Sonnet-20241022 58.93 60.00 61.93 69.55 55.95 47.22
Gemini-1.5-Pro-Latest 56.88 60.05 48.87 65.33 64.25 45.88
GLM-4v-Plus 54.09 50.00 48.71 57.20 67.50 47.05
Llama-3.2-90B-Vision-Instruct 51.40 57.60 54.47 55.03 56.50 33.43
GPT-4o-20241120 50.19 45.10 37.96 51.47 70.00 46.41
MiniCPM-V-2.6 48.72 48.04 42.80 44.55 59.75 48.44
Qwen2-VL-7B-Instruct 40.76 34.07 50.95 48.20 33.75 36.84
Llama-3.2-11B-Vision-Instruct 40.72 46.08 38.69 36.95 47.75 34.12
GPT-4o-Mini 37.44 37.26 37.29 35.86 44.00 32.79
Yi-Vision-V2 35.89 19.61 31.41 37.39 62.75 28.32
Table 8: Performance of LLMs on Values assessment across multiple scientific disciplines.
Discipline Overall Life Chemistry Earth Mathematics Physics
@@
The evaluation dataset is constructed following the methodology outlined in Section 3.2. The evaluated LLMs encompass both closed-source and open-source models, with a specific focus on their multimodal and non-multimodal capabilities. Specifically, our evaluation includes SOTA proprietary models such as GPT-4o-20241120 [81], GPT-4o-Mini [81], O1-Mini-20240912 [82], ERNIE-4.0-Turbo-8K-Latest [83], Claude-3.5-Sonnet-20241022 [84], Doubao-Pro-32K [85], Gemini-1.5-Pro-Latest [86], GLM-4-Plus [87], GLM-4v-Plus [88], Moonshot-V1-32K [89], Qwen-Plus [90], Yi-Lightning [91], and Yi-Vision-V2 [92]. Among these, Claude-3.5-Sonnet-20241022, Gemini-1.5-Pro-Latest, GLM-4v-Plus, GPT-4o-20241120, GPT-4o-Mini, and Yi-Vision-V2 incorporate multimodal capabilities, enabling them to process both textual and visual inputs. Additionally, we evaluated a broad spectrum of open-source models, highlighting their accessibility and customizability for AI4Science applications. Our evaluation includes DeepSeek-V3 [93], GLM-4-9B-Chat [94], InternLM2.5-20B-Chat [95], InternLM3-8B-Instruct [96], Llama3.1-8B-Instruct [97], Llama-3.2-11B-Vision-Instruct [98], Llama3.1-70B-Instruct [99], Llama-3.2-90B-Vision-Instruct [100], MiniCPM3-4B [101], MiniCPM-V-2.6 [102], Mistral-8B-Instruct-2410 [103], Mistral-Large-Instruct-2411 [104], Pixtral-Large-Instruct-2411 [105], Qwen2.5-72B-Instruct [106], Qwen2.5-7B-Instruct [107], Qwen2-VL-7B-Instruct [108], and Yi-34B-Chat [109]. Among these, Llama-3.2-11B-Vision-Instruct, Llama-3.2-90B-Vision-Instruct, MiniCPM-V-2.6, and Qwen2-VL-7B-Instruct exhibit multimodal capabilities, extending their usability beyond textual tasks.
4.2.1 Overall Performance
Table 3 presents the overall performance of each model across multiple scientific disciplines. Among all models, the Chinese open-source model DeepSeek-V3 achieves the highest overall score of 67.29, demonstrating a well-balanced capability across various knowledge domains. Claude-3.5-Sonnet (65.01) and O1-Mini (64.58) closely follow, highlighting their competitive performance. Additionally, the open-source model Llama3.1-70B and the closed-source model GLM-4-Plus also rank among the top-performing models. Among other domestic models, Qwen-Plus and Doubao-Pro-32K exhibit strong performance, reflecting their balanced and comprehensive capabilities.
From a disciplinary perspective, DeepSeek-V3 achieves the highest scores in Life Sciences (65.68) and Earth and Space Sciences (73.60), demonstrating strong domain knowledge in Life Sciences as well as spatial reasoning and map-based knowledge comprehension.O1-Mini excels in Mathematics (71.69) and Physics (63.56), demonstrating superior numerical and physical reasoning. However, its performance in other subjects is comparatively weaker. Meanwhile, Claude-3.5-Sonnet exhibits the strongest Chemistry understanding (73.05), outperforming all other models in this domain. These results highlight the diverse strengths of different LLMs, with DeepSeek-V3 excelling in Life and Earth sciences, O1-Mini in quantitative reasoning, and Claude-3.5-Sonnet in Chemistry, showcasing distinct model specializations in AI4Science applications.
To further analyze model performance across different evaluation dimensions, Figure 2 presents radar charts depicting the capabilities of the top-performing LLMs in five key dimensions: Knowledge, Understanding, Reasoning, Multimodality, and Values along with the overall performance. The visualization highlights how LLMs vary in strengths across disciplines and dimensions, providing deeper insights into their specialization. Notably, DeepSeek-V3 exhibits consistently high scores across most dimensions, particularly excelling in Understanding and Reasoning. Claude-3.5-Sonnet outperforms other models in Values and Chemistry-related tasks, reinforcing its strong domain-specific alignment. O1-Mini, despite leading in numerical disciplines such as Mathematics and Physics, shows comparatively lower scores in other areas. Additionally, multimodal models such as GPT-4o and Gemini-1.5-Pro demonstrate competitive performance in disciplines that require cross-modal information integration. These radar charts illustrate the varying trade-offs among LLMs, emphasizing their domain-specific capabilities and the importance of tailored model selection for different AI4Science applications.
4.2.2 Knowledge Evaluation
We evaluated the ability of LLM to acquire and apply key knowledge in different disciplines (Table 4). DeepSeek-V3 achieves the highest knowledge score (53.17), excelling in Life Sciences (58.61) and Mathematics (58.32). Qwen2.5-72B-Instruct (51.01) and Qwen-Plus (50.93) also perform well, suggesting that domestic open-source models can compete effectively in knowledge retention. Notably, Claude-3.5-Sonnet outperforms all models in Chemistry (61.43), highlighting its domain-specific factual accuracy in molecular sciences. DeepSeek-V3 leads in Physics (47.64), suggesting better recall of fundamental scientific principles.
4.2.3 Understanding Evaluation
We evaluated the ability of LLM to understand scientific facts (Table 5). DeepSeek-V3 (73.48) continues to lead, particularly excelling in Chemistry (85.00) and Mathematics (82.00). GPT-4o (70.27) ranks second overall, showing strong performance in Physics (69.00) and Life Sciences (62.33), highlighting its generalized reasoning ability. Claude-3.5-Sonnet (66.92) ranks third, achieving the highest Chemistry score (95.00), indicating superior contextual knowledge application in scientific disciplines. Interestingly, O1-Mini achieves the best performance in Mathematics (86.00), demonstrating strong numerical comprehension. These results suggest that while DeepSeek-V3 and GPT-4o offer balanced generalization, some models (e.g., Claude-3.5-Sonnet, O1-Mini) specialize in specific domains.
4.2.4 Reasoning Evaluation
For the Reasoning (Table 6), O1-Mini (85.25) outperforms all models, demonstrating SOTA reasoning capabilities, especially in Mathematics (87.09) and Physics (85.05). DeepSeek-V3 (79.66) and Gemini-1.5-Pro-Latest (79.07) also perform well, with DeepSeek-V3 excelling in Earth and Space Sciences (92.46). GPT-4o (72.36) and Claude-3.5-Sonnet (74.89) rank among the top models, confirming their robust logical inference abilities. These results highlight that reasoning ability varies significantly across models, with O1-Mini and DeepSeek-V3 excelling in mathematical and spatial logic, while Claude-3.5-Sonnet and GPT-4o demonstrate strong general reasoning across multiple subjects.
4.2.5 Multimodality Evaluation
For different LLMs’ abilities in understanding and reasoning about scientific charts (Table 7), Claude-3.5-Sonnet (58.93) ranks highest, particularly excelling in Chemistry (61.93) and Earth and Space Sciences (69.55). Gemini-1.5-Pro-Latest (56.88) and GLM-4v-Plus (54.09) also demonstrate strong multimodality capabilities. These results indicate that multimodality learning remains an evolving challenge, with closed-source models (e.g., Claude, Gemini) maintaining a significant advantage in integrating visual and textual reasoning.
4.2.6 Values Evaluation
For assessing whether different LLMs adhere to ethical guidelines and responsible AI principles (Table 8), Claude-3.5-Sonnet achieves the highest score (67.67), demonstrating the strongest alignment with ethical standards. Gemini-1.5-Pro-Latest (66.33) and GLM-4v-Plus (64.03) also perform well, reflecting their reinforced alignment mechanisms. Discipline Breakdown: Claude-3.5-Sonnet attains the highest scores in Chemistry (73.04) and Earth and Space Sciences (74.46). Gemini-1.5-Pro-Latest leads in Life Sciences (66.07) and Earth and Space Sciences (75.04), indicating strong alignment across the natural sciences. DeepSeek-V3 (62.86) maintains a balanced performance across multiple disciplines. These results suggest that commercial models tend to achieve superior value alignment, likely due to reinforced instruction tuning and safety fine-tuning techniques.
5 Conclusion
This study introduced SciHorizon, a comprehensive assessment framework for AI4Science readiness, integrating assessments of AI-ready scientific datasets and LLM capabilities. The framework systematically evaluated scientific datasets across four key dimensions to identify high-quality data for AI-driven research while assessing LLMs across five core indicators spanning major scientific disciplines. Based on data papers from 2018 to 2023, we provided dataset recommendations for Earth and Life Sciences. Meanwhile, by benchmarking more than 20 representative large models, we gained critical insights into their strengths and limitations. SciHorizon establishes a structured methodology for dataset curation and model assessment, contributing to AI-driven scientific discovery.

### v3

@@ B.2 More Details for LLM Assessment
Gemini-2.5-flash-preview 73.10 55.48 96.00 60.00 85.00 69.00
Claude-3.7-sonnet-thinking 71.19 51.37 97.00 59.60 78.00 70.00
Grok-3 70.51 57.53 96.00 61.00 74.00 64.00
@@
Gemini-2.0-Flash 69.22 52.74 93.94 58.00 75.76 65.66
Mistral-Medium-3 68.98 58.90 80.00 58.00 80.00 68.00
Claude-3.7-Sonnet 68.73 48.63 96.00 58.00 71.00 70.00
Qwen2.5-Max 67.92 54.79 83.00 59.00 79.80 63.00
@@
Table S8: Performance of LLMs on Reasoning assessment across multiple scientific disciplines.
@@
Gemini-2.5-pro-preview 92.10 87.55 93.55 96.23 92.31 90.84
Gemini-2.5-flash-preview-thinking 88.64 82.78 83.41 97.87 89.73 89.42
O3-Mini 88.16 83.07 80.68 98.41 87.44 91.22
Claude-3.7-sonnet-thinking 86.33 80.50 79.45 98.14 82.58 91.00
DeepSeek-R1 85.89 80.29 78.33 97.34 87.44 86.06
Qwen3-235B-A22B 85.35 82.73 79.16 97.61 82.03 85.24
@@
Gemini-2.0-Pro 82.40 81.64 76.74 98.14 73.65 81.84
QwQ-32B 82.32 76.28 76.74 97.08 83.42 78.09
Llama-4-Maverick-17B-128E 81.86 81.36 79.85 93.25 72.95 81.87
Qwen3-32B 80.63 75.43 71.21 98.14 81.69 76.69
Qwen2.5-Max 80.46 77.93 78.86 92.73 69.45 83.36
Llama-4-Scout-17B-16E 80.14 76.28 74.92 90.29 78.97 80.25
@@
Gemini-2.5-flash-preview 78.64 81.64 77.95 94.64 64.27 74.69
Mistral-Medium-3 76.39 73.43 73.71 93.25 63.73 77.85
Claude-3.7-Sonnet 76.27 74.07 74.92 94.90 54.44 83.02
Gemini-2.0-Flash 75.10 75.19 69.97 94.90 59.05 76.40
@@
Qwen3-30B-A3B 53.60 48.14 62.04 76.81 51.86 29.15
Table S9: Performance of LLMs on Values assessment across multiple scientific disciplines.
@@
Claude-3.7-sonnet-thinking 69.90 65.99 72.53 70.29 71.05 69.67
DeepSeek-R1 68.27 61.39 69.54 78.46 63.16 68.81
@@
Mistral-Small-24B-Instruct-2501 66.14 63.01 64.63 76.96 62.65 63.45
Claude-3.5-Haiku-20241022 66.03 62.33 66.83 69.52 63.68 67.78
O3-Mini 64.90 62.67 66.19 67.63 64.02 64.00
@@
Gemini-2.5-flash-preview 62.01 60.21 61.69 68.35 60.69 59.12
Gemini-2.5-flash-preview-thinking 61.80 60.54 63.64 70.10 60.85 53.88
@@
ERNIE-4.0-8K-Latest 41.57 37.84 43.98 51.18 40.55 34.31